Suppose the spot exchange rate for the Hungarian forint is HUF205.42. The inflation rate in the United States is 2.5 percent per year and is 5.5 percent in Hungary.

  

What do you predict the exchange rate will be in one year? (Do not round intermediate calculations and round your answer to 2 decimal places, e.g., 32.16.)

  

  Exchange rate HUF /$

  

What do you predict the exchange rate will be in two years? (Do not round intermediate calculations and round your answer to 2 decimal places, e.g., 32.16.)

  

  Exchange rate HUF /$

  

What do you predict the exchange rate will be in five years? (Do not round intermediate calculations and round your answer to 2 decimal places, e.g., 32.16.)

  

  Exchange rate HUF /$
Answer #1
Future rate=Spot rate*((1+Quoted currency Inflation rate)/(1+Base currency Inflation rate))^time
Future rate=205.42*((1+0.055)/(1+0.025))^1
Future rate(HUF/USD) = 211.43
Future rate=Spot rate*((1+Quoted currency Inflation rate)/(1+Base currency Inflation rate))^time
Future rate=205.42*((1+0.055)/(1+0.025))^2
Future rate(HUF/USD) = 217.62
Future rate=Spot rate*((1+Quoted currency Inflation rate)/(1+Base currency Inflation rate))^time
Future rate=205.42*((1+0.055)/(1+0.025))^5
Future rate(HUF/USD) = 237.29